A proton moving horizontally at 3.5×10⁴ m/s enters a uniform magnetic field of 0.075 T directed upward at 30° with the horizontal. Find the magnitude of the magnetic force acting on the proton

1
Expert's answer
2022-04-17T17:06:19-0400
"F=qvB\sin\theta\\
=1.6*10^{-19}*3.5*10^4*0.075*\sin30^\circ\\
=2.1*10^{-16}\:\rm N"

The force is directed out of the page.